OpenStudy (anonymous):

Can anyone plz explain me the main concept of mapping reducibility on Turing machine and what the symbols stands for plz

OpenStudy (kenljw):

If you mean the computer created by Turnig dot or no space AND, AB + for OR, A+B these are logic symbol used generally where A, B are binary bits Having to registers of binary bits one can add with binary addition using the above logic in proper sequence, as half adder or full adder

OpenStudy (kenljw):

AB 00=0 01=0 10=1 11=1 A+B 0+0=0 0+1=1 1+0=1 1+1=1